The sum of two numbers is four. The difference between ten and twice the smaller number is fourteen less than four times the larger.
Answer by mananth(16946)

x+y=4.................1
10-2x =4y-14
-2x-4y=-14-10
-2x-4y=-24.............2
multiply equation 1 by 2
2x+2y=8
add it to equation2
-2x+2x+2y-4y=-24+8
-2y=-16
y=8
so x=-4
...
m.ananth@hotmail.ca